# Knowledge-Driven Federated Graph Learning on Model Heterogeneity

### v3

@@ Abstract.
Federated graph learning (FGL) has emerged as a promising paradigm for collaborative graph representation learning, enabling multiple parties to jointly train models while preserving data privacy. However, most existing approaches assume homogeneous client models and largely overlook the challenge of model-centric heterogeneous FGL (MHtFGL), which frequently arises in practice when organizations employ graph neural networks (GNNs) of different scales and architectures. Such architectural diversity not only undermines smooth server-side aggregation, which presupposes a unified representation space shared across clients’ updates, but also further complicates the transfer and integration of structural knowledge across clients. To address this issue, we propose the Federated Graph Knowledge Collaboration (FedGKC) framework. FedGKC introduces a lightweight Copilot Model on each client to facilitate knowledge exchange while local architectures are heterogeneous across clients, and employs two complementary mechanisms: Client-side Self-Mutual Knowledge Distillation, which transfers effective knowledge between local and copilot models through bidirectional distillation with multi-view perturbation; and Server-side Knowledge-Aware Model Aggregation, which dynamically assigns aggregation weights based on knowledge provided by clients. Extensive experiments on eight benchmark datasets demonstrate that FedGKC achieves an average accuracy gain of 3.74% over baselines in MHtFGL scenarios, while maintaining excellent performance in homogeneous settings.
@@ 1. Introduction
In today’s data-driven world, graphs have become indispensable for representing complex relationships in domains such as social networks (Kumar et al., 2022), bioinformatics (Zhang et al., 2021b), and recommendations (He et al., 2020). Graphs effectively capture relational attributes and provide rich insights for machine learning. Graph Neural Networks (GNNs) (Wu et al., 2020) extend deep learning to non-Euclidean domains through message-passing mechanisms, enabling robust modeling that integrates local node features with global structural information. However, the rising data volume and privacy concerns expose the limitations of centralized training, which risks data leakage and hinders cross-entity collaboration. FGL (Liu et al., 2024a) addresses these issues by training models locally and sharing only model updates, thus enabling collaborative learning while guaranteeing privacy and communication efficiency.
@@
As FGL applications expand (Long et al., 2020), the disparity between clients in data volume, computational resources, and data quality becomes non-negligible. Larger enterprises typically possess abundant data, allowing them to deploy large-scale GNNs (Ji et al., 2021; Ying et al., 2018; Xu et al., 2018b). In contrast, smaller companies with limited data may adopt lightweight or domain-specific models tailored to their unique operational contexts (Veličković et al., 2018). Existing research in FGL has primarily addressed data-centric heterogeneity (Huang et al., 2023; Li et al., 2023b), while paying limited attention to the more fundamental model-centric heterogeneity challenge (MHtFGL), where divergent model architectures across clients introduce incompatibilities in representation learning and hinder effective knowledge integration at the server. Most FGL studies persuppose the homogeneous model architecture across clients, which curtails their applicability in practice. To bridge this gap, we propose a novel FGL framework designed to tackle the MHtFGL challenge while enhancing performance of collaborative training.
@@
From a model-centric perspective, the essence of FGL lies in the client–server paradigm, wherein locally learned implicit knowledge must be effectively communicated between clients and the server to facilitate collaborative learning. The key to effective collaboration under MHtFGL lies in the choice of a proxy Knowledge Carrier, an alternative medium that replaces the shared model parameters (Liang et al., 2020; Yi et al., 2023a), which have long served as the default knowledge carrier but require an identical local model architecture for parameter alignment and convergence. Prior works have proposed alternative knowledge carriers, which can be boardly summarized into three categories: (1) Prototype (Tan et al., 2022a; Zhang et al., 2024a; Yi et al., 2023b): They condense locally learned node representations into class-wise prototype vectors, which are typically low-dimensional embeddings summarizing local semantics. (2) Public/Synthetic Information (Fang and Ye, 2024; Yang et al., 2023; Zhang et al., 2024b): The server stores either the Public dataset or collaboratively trained synthetic Generators, which are utilized as the reference medium. They are distributed to local clients, who then align their local training using knowledge carriers such as prediction logits, gradients, or generator parameters, etc. (3)Partial Model (Wu et al., 2024; Alam et al., 2022; Lu et al., 2021): The server either distributes a homogeneous model component shared across heterogeneous clients or allocates model subsets of varying depth or parameter scale according to each client’s computational capacity and the relevance of parameter subsets to its local data distribution. Within the context of FGL, Parial Model can be regarded as the most effective solution since the strength of GNNs stems from the neighbor aggregation and propagation mechanism, through which implicit graph knowledge is tightly entangled with the trained model parameters themselves. In contrast, the knowledge carriers employed by the other two strategies rely primarily on model outputs (e.g., logits or gradients), which can only serve as indirect reflections of the underlying graph-structural knowledge.
@@
Building on the above analysis, we propose the Federated Graph Knowledge Collaboration (FedGKC) framework to address MHtFGL. The framework leverages knowledge distillation while explicitly adapting it to structural properties of graph data. To this end, we introduce copilot models on the client-side to bridge architectural differences and design a knowledge-aware aggregation strategy on the server-side to ensure effective integration of diverse knowledge uploaded from clients. Specifically, FedGKC consists of two key components: (1) Self-Mutual Knowledge Distillation (SMKD) on heterogeneous clients. We establish a bidirectional knowledge distillation process between each local odel and its copilot model. This design enables the copilot to acquire localized knowledge while simultaneously injecting global knowledge back into the local model. Beyond class-level predictions, SMKD also transfers neighborhood information, thereby enhancing the model’s ability to capture topological dependencies. (Sec. 3.3.1) To improve robustness, we incorporate a multi-view data perturbation strategy for local self-distillation , ensuring consistent predictions across perturbed local graphs and boosting generalizabilities. (Sec. 3.3.2) (2) Knowledge-Aware Model Aggregation (KAMA) on the server side. Unlike the conventional FedAvg algorithm (McMahan et al., 2017), which assigns weights to clients solely based on their data volume, KAMA recognizes that heterogeneous clients differ significantly in learning capacity and knowledge quality. Specifically, we design an adaptive weighting mechanism based on data volume, knowledge strength, and knowledge clarity. Performing aggregation on knowledge-aware weights, KAMA achieves more accurate and efficient integration between clients, guiding the global model toward better convergence.
To summarize, our work offers three following contributions: (1) Problem Identification. We are the first to systematically investigate the challenge of MHtFGL, specifically the partial model strategy, where diverse client architectures hinder unified training and complicate the integration of graph knowledge in real-world deployments. (2) Innovative Framework. We propose FedGKC, a knowledge-driven FGL framework that enables effective collaboration across clients with heterogeneous models. It incorporates two key mechanisms: SMKD, which bridges structural differences through bidirectional and topology-aware knowledge distillation between local and copilot models; and KAMA, which adaptively aggregates client contributions based on knowledge quality, ensuring more reliable global insights. (3) Superior Performance. We validate FedGKC on eight benchmark datasets under both heterogeneous and homogeneous settings. Experiments show consistent improvements, with an average accuracy gain of 3.74% and up to 8.13% over state-of-the-art baselines, confirming both the effectiveness and generalizability of our approach.
@@ 2.1. Federated Graph Learning
FGL extends FL to GNNs and enables collaborative training on graph-structured data while preserving data privacy. Instead of exchanging raw data, clients upload learned graph representation models to the secured server to aggregate cross-client knowledge. Existing FGL research can be broadly categorized into two settings: (1) Graph-FL: Each client owns multiple independent graphs and the goal is to address graph-level tasks (e.g., graph classification). Representative approaches include GCFL (Xie et al., 2021) and FedStar (Tan et al., 2023),etc. (2) Subgraph-FL: Each client holds a subgraph of an implicit global graph and the objective is to solve node-level tasks (e.g., node classification). Representative works include FGSSL (Huang et al., 2023), FedGTA (Li et al., 2023b), FedPub (Baek et al., 2023), and FedSage+ (Zhang et al., 2021c). Notably, previous studies have concentrated on addressing data heterogeneity challenges  (Huang et al., 2023; Li et al., 2023b). However, in real-world scenarios, GNNs from different clients may deploy divergent architectural models, posing fundamental obstacles to unified graph learning and complicating collaborative training. To the best of our knowledge, we are the first to investigate this issue and propose the new paradigm tailored for this challenge.
@@ 2.2. Model-Centric Heterogeneous FL
Model-centric heterogeneous federated learning (MHtFL) enables flexible and diverse client architectures while preserving data privacy. Existing MHtFL methods can be broadly classified into three categories, each with distinct limitations: (1) Prototype (Tan et al., 2022a; Zhang et al., 2024a; Tan et al., 2022b) avoid sharing parameters but suffer from biased classifiers, further exacerbated by architectural diversity (Li et al., 2023a), which undermines efficacy of the global knowledge aggregation. (2) Public/Synthetic Information (Fang and Ye, 2024; Yang et al., 2023; Zhang et al., 2024b, 2021a; Lin et al., 2020) distribute globally stored public dataset or generators for syntethizing graph information to clients. During local training, heterogeneous clients align their predictions or gradients on shared public data or upload generator parameters for aggregation. However, such proxies act as indirect knowledge carriers and fail to capture the topology-dependent structural knowledge that is crucial for effective FGL training. (3) Partial Model methods such as LG-FedAvg (Liang et al., 2020), FedGen (Zhu et al., 2021), and FedGH (Yi et al., 2023a) permit heterogeneity in major model components while enforcing homogeneity on a small shared module. This design enables limited global knowledge transfer but constrains model expressiveness. Approaches such as FML (Shen et al., 2023) and FedKD (Wu et al., 2022) employ auxiliary global models for mutual distillation; however, early training rounds risk exchanging low-quality or unstable knowledge.
@@ 2.3. Knowledge Distillation on Graph
Knowledge distillation (KD) (Hinton, 2015) transfers knowledge from large models to smaller ones, reducing computational cost while preserving performance. In graph learning, KD enables knowledge transfer across diverse models, with the main challenge being what knowledge to distill and how to distill it.
Three main categories of transferable knowledge are commonly explored: (1) Logits, i.e., outputs before the Softmax layer, where methods minimize distribution divergence to transfer predictive knowledge (Wang et al., 2021; Yan et al., 2020; Yang et al., 2021). (2) Graph structure, which captures node-edge connectivity and relationships, with approaches aligning nodes via neighborhood relations to transfer structural information (Yang et al., 2020; Guo et al., 2022). (3) Embeddings, i.e., intermediate node representations extracted from hidden layers, which contain feature and structural context. It can be distilled to guide another model’s training (Huo et al., 2023; He et al., 2022; Yu et al., 2022).
Regarding distillation methods, direct distillation minimizes differences between models, promoting intuitive knowledge transfer (Yan et al., 2020); adaptive distillation offers flexibility by considering the importance of different knowledge components. Methods like RDD (Zhang et al., 2020), FreeKD (Feng et al., 2022), and MulDE (Wang et al., 2021) adjust the distillation process based on the certainty or correctness of the knowledge, allowing models to focus more on reliable information while avoiding knowledge that is less informative or noisy. Furthermore, various distillation strategies tailored for specific tasks exist within machine learning (He et al., 2022). Our study aims to ensure the preservation and transfer of effective graph information within heterogeneous federated scenarios, thereby facilitating collaborative training among multiple clients.
@@ 3.3.1. Mutual distillation addressing model heterogeneity
In traditional knowledge distillation methods (Yang et al., 2021), knowledge is typically transferred by aligning the prediction results of different models on the same data. However, merely aligning the prediction results may limit the knowledge learned by the models, resulting in a lack of comprehensiveness and representativeness. In the context of graph data processed by GNNs, node representations are primarily derived from capturing neighborhood information. Therefore, we propose propagating neighborhood knowledge to enhance the utilization of topological information captured by GNNs:
@@
where $z$ and $h$ represent the embedding outputs of the feature extractor of the two models, respectively. $\mathcal{V}_{i}$ represents the set of neighbor nodes of node $i$, $\mathcal{N}$ denotes the total number of nodes, $\sigma=softmax(\cdot)$ denotes an activation function, and $\mathcal{D}(\cdot)$ denotes a distance function (e.g., Kullback Leibler divergence (Hinton, 2015)).
@@ 3.3.2. Self-distillation mining model potential
To further stimulate the performance of the local model, we introduce a self-distillation strategy. We first create two perspectives on the same graph data, a strongly perturbated view and a weakly perturbated view. This dual-perspective strategy effectively enhances the diversity of graph data by applying varying degrees of edge removal and feature masking (Zhao et al., 2022). The weakly perturbated view retains more original information, while the strongly perturbated view significantly obscures and removes more graph information (e.g., topological structure, node semantics). We expect the predictions of the strongly perturbated view to closely align with the predictions of the weakly perturbated view for the same input. It aims to ensure that, even when faced with substantial information alteration, the local model maintains consistency in its predictions, thereby deepening its understanding of the data’s intrinsic nature. The optimization objective for the self-distillation of the local model is denoted as

### v4

@@ Abstract.
Federated graph learning (FGL) has emerged as a promising paradigm for collaborative graph representation learning, enabling multiple parties to jointly train models while preserving data privacy. However, most existing approaches assume homogeneous client models and largely overlook the challenge of model-centric heterogeneous FGL (MHtFGL), which frequently arises in practice when organizations employ graph neural networks (GNNs) of different scales and architectures. Such architectural diversity not only undermines smooth server-side aggregation, which presupposes a unified representation space shared across clients’ updates, but also further complicates the transfer and integration of structural knowledge across clients. To address this issue, we propose the Federated Graph Knowledge Collaboration (FedGKC) framework. FedGKC introduces a lightweight Copilot Model on each client to facilitate knowledge exchange while local architectures are heterogeneous across clients, and employs two complementary mechanisms: Client-side Self-Mutual Knowledge Distillation, which transfers effective knowledge between local and copilot models through bidirectional distillation with multi-view perturbation; and Server-side Knowledge-Aware Model Aggregation, which dynamically assigns aggregation weights based on knowledge provided by clients. Extensive experiments on eight benchmark datasets demonstrate that FedGKC achieves an average accuracy gain of 3.88% over baselines in MHtFGL scenarios, while maintaining excellent performance in homogeneous settings.
@@ 1. Introduction
In today’s data-driven world, graphs have become indispensable for representing complex relationships in domains such as social networks (kumar2022socialgraph), bioinformatics (zhang2021biograph), and recommendations (he2020recomgraph). Graphs effectively capture relational attributes and provide rich insights for machine learning. Graph Neural Networks (GNNs) (wu2020GNN) extend deep learning to non-Euclidean domains through message-passing mechanisms, enabling robust modeling that integrates local node features with global structural information. However, the rising data volume and privacy concerns expose the limitations of centralized training, which risks data leakage and hinders cross-entity collaboration. FGL (liu2024FGL) addresses these issues by training models locally and sharing only model updates, thus enabling collaborative learning while guaranteeing privacy and communication efficiency.
@@
As FGL applications expand (long2020federatedProblem), the disparity between clients in data volume, computational resources, and data quality becomes non-negligible. Larger enterprises typically possess abundant data, allowing them to deploy large-scale GNNs (LargeCK; WebScale; jknet). In contrast, smaller companies with limited data may adopt lightweight or domain-specific models tailored to their unique operational contexts (velivckovic2018gat). Existing research in FGL has primarily addressed data-centric heterogeneity (ijcai2023FGSSL; li2023fedgta), while paying limited attention to the more fundamental model-centric heterogeneity challenge (MHtFGL), where divergent model architectures across clients introduce incompatibilities in representation learning and hinder effective knowledge integration at the server. Most FGL studies persuppose the homogeneous model architecture across clients, which curtails their applicability in practice. To bridge this gap, we propose a novel FGL framework designed to tackle the MHtFGL challenge while enhancing performance of collaborative training.
@@
From a model-centric perspective, the essence of FGL lies in the client–server paradigm, wherein locally learned implicit knowledge must be effectively communicated between clients and the server to facilitate collaborative learning. The key to effective collaboration under MHtFGL lies in the choice of a proxy Knowledge Carrier, an alternative medium that replaces the shared model parameters (liang2020LG; yi2023fedgh), which have long served as the default knowledge carrier but require an identical local model architecture for parameter alignment and convergence. Prior works have proposed alternative knowledge carriers, which can be boardly summarized into three categories: (1) Prototype (tan2022fedproto; zhang2024fedtgp; yi2023fedssa): They condense locally learned node representations into class-wise prototype vectors, which are typically low-dimensional embeddings summarizing local semantics. (2) Public/Synthetic Information (fang2024noise; yang2023allosteric; zhang2024improving): The server stores either the Public dataset or collaboratively trained synthetic Generators, which are utilized as the reference medium. They are distributed to local clients, who then align their local training using knowledge carriers such as prediction logits, gradients, or generator parameters, etc. (3)Partial Model (wu2024fiarse; alam2022fedrolex; lu2021heterogeneous): The server either distributes a homogeneous model component shared across heterogeneous clients or allocates model subsets of varying depth or parameter scale according to each client’s computational capacity and the relevance of parameter subsets to its local data distribution. Within the context of FGL, Parial Model can be regarded as the most effective solution since the strength of GNNs stems from the neighbor aggregation and propagation mechanism, through which implicit graph knowledge is tightly entangled with the trained model parameters themselves. In contrast, the knowledge carriers employed by the other two strategies rely primarily on model outputs (e.g., logits or gradients), which can only serve as indirect reflections of the underlying graph-structural knowledge.
@@
Building on the above analysis, we propose the Federated Graph Knowledge Collaboration (FedGKC) framework to address MHtFGL. The framework leverages knowledge distillation while explicitly adapting it to structural properties of graph data. To this end, we introduce copilot models on the client-side to bridge architectural differences and design a knowledge-aware aggregation strategy on the server-side to ensure effective integration of diverse knowledge uploaded from clients. Specifically, FedGKC consists of two key components: (1) Self-Mutual Knowledge Distillation (SMKD) on heterogeneous clients. We establish a bidirectional knowledge distillation process between each local odel and its copilot model. This design enables the copilot to acquire localized knowledge while simultaneously injecting global knowledge back into the local model. Beyond class-level predictions, SMKD also transfers neighborhood information, thereby enhancing the model’s ability to capture topological dependencies. (Sec. 3.3.1) To improve robustness, we incorporate a multi-view data perturbation strategy for local self-distillation , ensuring consistent predictions across perturbed local graphs and boosting generalizabilities. (Sec. 3.3.2) (2) Knowledge-Aware Model Aggregation (KAMA) on the server side. Unlike the conventional FedAvg algorithm (2017FedAvg), which assigns weights to clients solely based on their data volume, KAMA recognizes that heterogeneous clients differ significantly in learning capacity and knowledge quality. Specifically, we design an adaptive weighting mechanism based on data volume, knowledge strength, and knowledge clarity. Performing aggregation on knowledge-aware weights, KAMA achieves more accurate and efficient integration between clients, guiding the global model toward better convergence.
To summarize, our work offers three following contributions: (1) Problem Identification. We are the first to systematically investigate the challenge of MHtFGL, specifically the partial model strategy, where diverse client architectures hinder unified training and complicate the integration of graph knowledge in real-world deployments. (2) Innovative Framework. We propose FedGKC, a knowledge-driven FGL framework that enables effective collaboration across clients with heterogeneous models. It incorporates two key mechanisms: SMKD, which bridges structural differences through bidirectional and topology-aware knowledge distillation between local and copilot models; and KAMA, which adaptively aggregates client contributions based on knowledge quality, ensuring more reliable global insights. (3) Superior Performance. We validate FedGKC on eight benchmark datasets under both heterogeneous and homogeneous settings. Experiments show consistent improvements, with an average accuracy gain of 3.88% and up to 8.13% over state-of-the-art baselines, confirming both the effectiveness and generalizability of our approach.
@@ 2.1. Federated Graph Learning
FGL extends FL to GNNs and enables collaborative training on graph-structured data while preserving data privacy. Instead of exchanging raw data, clients upload learned graph representation models to the secured server to aggregate cross-client knowledge. Existing FGL research can be broadly categorized into two settings: (1) Graph-FL: Each client owns multiple independent graphs and the goal is to address graph-level tasks (e.g., graph classification). Representative approaches include GCFL (xie2021GCFL) and FedStar (tan2023fedstar),etc. (2) Subgraph-FL: Each client holds a subgraph of an implicit global graph and the objective is to solve node-level tasks (e.g., node classification). Representative works include FGSSL (ijcai2023FGSSL), FedGTA (li2023fedgta), FedPub (baek2023fedpub), and FedSage+ (zhang2021fedsage). Notably, previous studies have concentrated on addressing data heterogeneity challenges  (ijcai2023FGSSL; li2023fedgta). However, in real-world scenarios, GNNs from different clients may deploy divergent architectural models, posing fundamental obstacles to unified graph learning and complicating collaborative training. To the best of our knowledge, we are the first to investigate this issue and propose the new paradigm tailored for this challenge.
@@ 2.2. Model-Centric Heterogeneous FL
Model-centric heterogeneous federated learning (MHtFL) enables flexible and diverse client architectures while preserving data privacy. Existing MHtFL methods can be broadly classified into three categories, each with distinct limitations: (1) Prototype (tan2022fedproto; zhang2024fedtgp; tan2022fedpcl) avoid sharing parameters but suffer from biased classifiers, further exacerbated by architectural diversity (li2023bias), which undermines efficacy of the global knowledge aggregation. (2) Public/Synthetic Information (fang2024noise; yang2023allosteric; zhang2024improving; zhang2021parameterized; lin2020ensemble) distribute globally stored public dataset or generators for syntethizing graph information to clients. During local training, heterogeneous clients align their predictions or gradients on shared public data or upload generator parameters for aggregation. However, such proxies act as indirect knowledge carriers and fail to capture the topology-dependent structural knowledge that is crucial for effective FGL training. (3) Partial Model methods such as LG-FedAvg (liang2020LG), FedGen (zhu2021fedgen), and FedGH (yi2023fedgh) permit heterogeneity in major model components while enforcing homogeneity on a small shared module. This design enables limited global knowledge transfer but constrains model expressiveness. Approaches such as FML (shen2023fedfml) and FedKD (wu2022communication) employ auxiliary global models for mutual distillation; however, early training rounds risk exchanging low-quality or unstable knowledge.
@@ 2.3. Knowledge Distillation on Graph
Knowledge distillation (KD) (hinton2015distilling) transfers knowledge from large models to smaller ones, reducing computational cost while preserving performance. In graph learning, KD enables knowledge transfer across diverse models, with the main challenge being what knowledge to distill and how to distill it.
Three main categories of transferable knowledge are commonly explored: (1) Logits, i.e., outputs before the Softmax layer, where methods minimize distribution divergence to transfer predictive knowledge (wang2021MulDElogitsKD1; yan2020logitsKD2; yang2021logitsKD3). (2) Graph structure, which captures node-edge connectivity and relationships, with approaches aligning nodes via neighborhood relations to transfer structural information (yang2020structKD1; guo2022structKD2). (3) Embeddings, i.e., intermediate node representations extracted from hidden layers, which contain feature and structural context. It can be distilled to guide another model’s training (huo2023embeddingKD1; he2022embeddingKD2; yu2022embeddingKD3).
Regarding distillation methods, direct distillation minimizes differences between models, promoting intuitive knowledge transfer (yan2020logitsKD2); adaptive distillation offers flexibility by considering the importance of different knowledge components. Methods like RDD (zhang2020RDD), FreeKD (feng2022freekd), and MulDE (wang2021MulDElogitsKD1) adjust the distillation process based on the certainty or correctness of the knowledge, allowing models to focus more on reliable information while avoiding knowledge that is less informative or noisy. Furthermore, various distillation strategies tailored for specific tasks exist within machine learning (he2022embeddingKD2). Our study aims to ensure the preservation and transfer of effective graph information within heterogeneous federated scenarios, thereby facilitating collaborative training among multiple clients.
@@ 3.3.1. Mutual distillation addressing model heterogeneity
In traditional knowledge distillation methods (yang2021logitsKD3), knowledge is typically transferred by aligning the prediction results of different models on the same data. However, merely aligning the prediction results may limit the knowledge learned by the models, resulting in a lack of comprehensiveness and representativeness. In the context of graph data processed by GNNs, node representations are primarily derived from capturing neighborhood information. Therefore, we propose propagating neighborhood knowledge to enhance the utilization of topological information captured by GNNs:
@@
where $z$ and $h$ represent the embedding outputs of the feature extractor of the two models, respectively. $\mathcal{V}_{i}$ represents the set of neighbor nodes of node $i$, $\mathcal{N}$ denotes the total number of nodes, $\sigma=softmax(\cdot)$ denotes an activation function, and $\mathcal{D}(\cdot)$ denotes a distance function (e.g., Kullback Leibler divergence (hinton2015distilling)).
@@ 3.3.2. Self-distillation mining model potential
To further stimulate the performance of the local model, we introduce a self-distillation strategy. We first create two perspectives on the same graph data, a strongly perturbated view and a weakly perturbated view. This dual-perspective strategy effectively enhances the diversity of graph data by applying varying degrees of edge removal and feature masking (zhao2022graphAug). The weakly perturbated view retains more original information, while the strongly perturbated view significantly obscures and removes more graph information (e.g., topological structure, node semantics). We expect the predictions of the strongly perturbated view to closely align with the predictions of the weakly perturbated view for the same input. It aims to ensure that, even when faced with substantial information alteration, the local model maintains consistency in its predictions, thereby deepening its understanding of the data’s intrinsic nature. The optimization objective for the self-distillation of the local model is denoted as